OpenStudy (anonymous):

2. How many solutions does this system have? 2x + y = 3 6x = 9 – 3y (1 point)

OpenStudy (hba):

Well,you can solve it and check. There are a number of methods you can use 1)Elimination 2)Substitution

OpenStudy (anonymous):

I was not in class for this whole past week. I have been very ill. I dont know what any of that means.

OpenStudy (hba):

Well lets see it like this 2x+y=3---(1) 6x+3y=9--(2) Multiply Eq(1) by 2. And, then subract these equations.

OpenStudy (anonymous):

Okay, I am still kind of confused. Im sorry. /:

OpenStudy (hba):

Did you Multiply Eq(1) by 2,What did you get.

OpenStudy (anonymous):

No, hold on I will do that now.

OpenStudy (hba):

ok.

OpenStudy (hba):

Take your time...

OpenStudy (anonymous):

I am lost. I am horrible at math. ...

OpenStudy (hba):

What did you get ?

OpenStudy (anonymous):

Nothing. . . /':

OpenStudy (anonymous):

I can't do it. It is all giberish to me...

OpenStudy (hba):

So, you are doing something you dont know. Lets start working on your basics then or Lets use the other method what do you think ?

OpenStudy (anonymous):

Yes, and it doesnt matter.

OpenStudy (anonymous):

X=3/2 and y=0

OpenStudy (hba):

@chow Giving direct answers is not allowed Please check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (hba):

2x + y = 3--(1) 6x = 9 – 3y--(2) (1)=) y=3-2x---(3) Now put Eq(3) or y in Eq(2)

OpenStudy (anonymous):

That wasnt a direct answer... The answer will be either, 1, none, 2 or infinite. So @chow is okay..

OpenStudy (anonymous):

thanks Kierstan...I don't have energy to type total derivation..

OpenStudy (anonymous):

You're welcome.

OpenStudy (hba):

@KierstanDanielle If you have the options then you can actually try putting your answer options and checking if L.H.S=R.H.S If L.H.S=R.H.S then that is the correct answer option.

OpenStudy (hba):

@chow I hope you understood what i said.

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

What does L.H.S. and R.H.S mean?

OpenStudy (hba):

Left hand side = Right hand side

OpenStudy (anonymous):

oh. hahha

OpenStudy (hba):

@chow Thanks a lot for understanding.

OpenStudy (anonymous):

Okay, I think I might have the answer. Thanks for helping !

OpenStudy (hba):

Your'e welcome.